# Bayesian Estimation and Tuning-Free Rank Detection for Probability Mass Function Tensors

### v1

@@ V-A1 VB-PMF performance
Figs. 2 and 3 presents results (averaged over 100 independent trials) for a rank-$5$ and a rank-$10$ PMF tensor, respectively. For both outage values $p=\{0,0.3\}$, the rank inferred by VB-PMF converges to the true rank as the number of observations is increased. A similar observation can be made for BIC and DNML. However, for $R=10$ and $p=0.3$, BIC and DNML are not able to detect the correct rank and would need much more data to do so. On the other hand, AIC exhibits a large variance in the inferred rank, which is unsurprising as AIC is not a consistent criterion [31]. In terms of the accuracy, VB-PMF performs comparably to or better than the other model order estimation techniques. Here, the ‘Oracle’ ML estimate (MLE), which assumes that the true hidden state associated with each observation is known, serves as an empirical lower bound for the KLD. This result can be attributed to the fact that VB-PMF detects the correct rank with minimal or no misfit as the size of the dataset increases. Moreover, the runtime plots show that VB-PMF provides significant gains in computational efficiency as the size of the dataset grows, even though the runtime increases linearly with the dataset size. This is because the rank is estimated via a single training run, while the other model order selection techniques need multiple training runs over the candidate ranks.
Fig. 4 shows the rank estimation performance as a function of the loading vector hyperparameter $\alpha_{\lambda}$. The estimated rank is consistent regardless of the choice of $\alpha_{\lambda}$. It is only required that $\alpha_{\lambda}\ll 1$ in order to obtain a sparse loading vector for rank inference. As mentioned earlier, in the absence of prior information, it is reasonable to set $\alpha_{n,r}=1$ in order to allow all possible solutions for the factor matrix columns. This result demonstrates the robustness of VB-PMF to the choice of the hyperparameter $\alpha_{\lambda}$.
V-A2 SVB-PMF performance
TABLE I: Performance of different algorithms on the MovieLens dataset.
Estimation
algorithm
Model order
selection criterion
RMSE MAE NLL
Most common
CPD rank
Mean runtime (hrs)
VB-PMF (proposed) - $\mathbf{0.726\pm 0.005}$ $\mathbf{0.540\pm 0.003}$ $\mathbf{20.543\pm 0.080}$ $44$ $3.17$
SVB-PMF (proposed) - $0.737\pm 0.006$ $0.551\pm 0.004$ $20.594\pm 0.082$ $31$ $\mathbf{0.52}$
SQUAREM-PMF [12] ValErr $0.731\pm 0.005$ $0.544\pm 0.003$ $20.552\pm 0.083$ $30$ $21.27$
AIC [13] $0.732\pm 0.006$ $0.544\pm 0.004$ $20.553\pm 0.079$ $30$ $59.53$
BIC [14] $0.790\pm 0.010$ $0.592\pm 0.008$ $20.727\pm 0.083$ $10$ $59.53$
DNML [15] $0.759\pm 0.010$ $0.567\pm 0.008$ $20.638\pm 0.088$ $15$ $59.98$
CTF3D [5] ValErr $0.846\pm 0.020$ $0.617\pm 0.022$ $21.838\pm 0.139$ $30$ $41.55$
BMF [32] ValErr $0.958\pm 0.006$ $0.740\pm 0.004$
Next, we consider a larger problem with $N=50$ discrete random variables each taking $I_{n}=10$ states and with rank $R=25$. The outage probability is set to $p=0.3$. As before, the elements of $\bm{\lambda}$ and $\{\bm{A}_{n}\}_{n=1}^{N}$ are drawn from $U(0.3,1)$ and $U(0,1)$, respectively, and normalized to fulfill the probability simplex constraints.
We investigate three choices for the SVB-PMF minibatch size $M$. In particular, since we would like the asymptotic complexity of SVB-PMF to be sublinear in $T$, we set $M=T^{1/b}$, where $b\in\{1.5,2,3\}$. The performance of SVB-PMF is compared with that of VB-PMF. For both algorithms, the hyperparameters are set to $\alpha_{\lambda}=10^{-6}$ and $\alpha_{n,r}=1$, the pruning threshold is set to $\epsilon=10^{-3}$, and the initial rank is $R=50$. Furthermore, $10\%$ of the observations are held out as a test dataset to check for convergence. The results are averaged over 100 independent trials.
In Fig. 5, we present the mean estimated rank as a function of the number of observations $T$ in the dataset. For $T=10^{4}$, all algorithms are unable to determine the correct model order due to an insufficient number of samples. As $T$ is increased, we observe that VB-PMF accurately determines the model order. On the other hand, while SVB-PMF benefits from larger minibatch sizes for intermediate values of $T$, overfitting is observed for $T=\{5\cdot 10^{5},{\color[rgb]{0,0,0}10^{6}}\}$, although to a reduced extent for the smallest minibatch size, $M=T^{1/3}$. The overfitting in our experiments arises as a result of imposing a maximum allowed runtime for both algorithms, as well as a convergence threshold. Higher accuracy can be obtained by increasing the maximum allowed runtime and/or decreasing the convergence threshold. The plot on the right in Fig. 5 shows a reduction in overfitting for both algorithms when the maximum allowed runtime is increased from 10 hours to 15 hours with the convergence threshold set to zero.
To gain further insight, we study the mean negative log-likelihood (NLL) of the held-out test dataset, which serves as an indicator of the quality of the learned model. As shown in Fig. 7, the mean NLL for all three SVB-PMF variants approaches the ground truth as $T$ increases, with larger minibatch sizes leading to faster convergence. Even though there is some overfitting for $T=\{5\cdot 10^{5},10^{6}\}$ (see Fig. 5), the effect on the quality of the model is quite minimal.
Finally, in Fig. 7, we present the distribution of the runtime for each value of $T$. For $T=10^{4}$, both VB-PMF and SVB-PMF have roughly equivalent runtimes, while VB-PMF is more computationally efficient for $T=5\cdot 10^{4}$. For moderately large values of $T$, VB-PMF is able to converge faster by analyzing the entire dataset, while SVB-PMF takes longer due to the minibatch operation. However, for $T=\{5\cdot 10^{5},10^{6}\}$, the linear dependence of VB-PMF on $T$ starts to become a bottleneck and the computational advantage of SVB-PMF is readily apparent. Note that even though one could argue that such a large dataset is not required in this case (cf. Fig. 7), in a real-world scenario, the true PMF (and, hence, the true NLL) is unknown. Therefore, it would be difficult to know in advance how much data is required for a good estimate.
V-B Real-Data Experiments
@@
We evaluate the performance of the proposed algorithms in a real-world movie recommendation application. For this purpose, we utilize the MovieLens 25M dataset [33] which contains 25 million movie ratings on a 5-star scale with half-star increments, i.e., $\{0.5,\dotsc,5.0\}$. From the entire dataset, we select the ratings of $N=50$ top-rated movies. We perform some pre-processing by omitting users who have rated only one or none of the selected movies and mapping the ratings to a discrete scale $\{1,\dotsc,10\}$. Thus, we obtain a movie-user rating matrix $\bm{Y}\in\mathbb{R}^{N\times T}$ where the number of users $T=1.{\color[rgb]{0,0,0}5}2\times 10^{5}$. The percentage of missing ratings in the rating matrix is about 67% (outage probability $p=0.67$). Moreover, since we are dealing with real data, the true rank (model order) is unknown.
@@
Our aim is to estimate the joint PMF (i.e., $\{\bm{A}_{n}\}_{n=1}^{N}$ and $\bm{\lambda}$) of the ratings from the observations in order to predict missing ratings. In this scenario, the joint PMF to be estimated is a 50-dimensional tensor $\bm{\mathcal{X}}\in\mathbb{R}^{I_{1}\times\cdots\times I_{N}}$ with each dimension $I_{n}=10$, $n=1,\dotsc,50$. We run 20 Monte Carlo trials whereby 70% of the ratings are used to estimate the joint PMF, 10% serve as a validation set for model order selection (choosing the CPD rank) or to check for convergence, and 20% are used for testing the algorithms. In each trial, a different training/validation/testing split is applied to the rating matrix.
The proposed algorithms, VB-PMF and SVB-PMF, require no cross-validation for model order selection. Hence, in both cases, the NLL of the validation set is used to test for convergence. Both algorithms are initialized with $R=50$ components and the hyperparameters $\alpha_{\lambda}$ and $\alpha_{n,r}$ are set to $10^{-6}$ and $1$, respectively. The batch size for SVB-PMF is set to $M=500$ ($\approx T^{1/2}$). For comparison, we include AIC, BIC, and DNML, for which the best model order is chosen based on the respective criterion evaluated on the training set. We also include ValErr, for which the best model order is chosen based on the root mean square error (RMSE) of the validation set, as is the standard practice in the machine learning literature. The estimation algorithm for these four techniques is SQUAREM-PMF [12] (cf. Section V-A). We also include coupled tensor factorization based on 3D marginals (CTFD) [5] as a benchmark for PMF estimation algorithms. For computational efficiency, the set of candidate ranks was restricted to $\{5,10,15,20,25,30\}$. These PMF-based approaches are compared with biased matrix factorization (BMF) [32] which is a classical technique for recommendation systems.
@@
During the testing phase (also during validation for the ValErr approach), for each test sample, one rating is hidden and predicted using the estimated PMF from each approach. The predicted rating is found by computing its conditional expectation with respect to the estimated joint PMF. In particular, let the current test sample be $\bm{y}_{t}=[y_{1,t},\dotsc,y_{N,t}]^{\mathsf{T}}$ and suppose that $y_{N,t}=0$, i.e., hidden. The predicted rating is then
$\hat{y}_{N,t}=\sum_{i_{N}=1}^{I_{N}}i_{N}\mathsf{Pr}(i_{N}\,|\,y_{1,t},\dotsc,y_{N-1,t}),$ (55)
where the conditional PMF can be computed from the joint PMF via Bayes’ theorem. The quality of the predictions is evaluated by computing the RMSE and the mean absolute error (MAE) between the predicted and true ratings. Moreover, the mean NLL of the test dataset is presented as a means of comparing the overall quality of the models (the lower the better).
Table I presents the performance of the various PMF estimation algorithms in terms of the RMSE, the MAE and the mean NLL. In addition, we also report the most common (estimated) rank as well as the mean runtime. It can be observed that VB-PMF achieves the best performance across all metrics. In comparison, SVB-PMF experiences some performance loss. This result is expected since SVB-PMF operates on minibatches of data while VB-PMF processes the entire dataset. However, SVB-PMF yields remarkable savings in the computation time compared to VB-PMF, and thus the marginal degradation in accuracy can be tolerated. As expected, there is a huge computational cost incurred by SQUAREM-PMF and CTF3D due to model order selection, although SQUAREM-PMF performs comparably to the Bayesian algorithms in some cases. Note that the number of candidate ranks for SQUAREM-PMF and CTF3D was restricted for computational efficiency; our approaches would be even more advantageous if more candidate ranks were considered. Compared to the other PMF estimation algorithms, CTF3D experiences performance degradation due to the restriction to only use 3D marginals. In terms of the RMSE and the MAE, all PMF estimation algorithms outperform BMF. This result demonstrates that the estimated joint PMF tensor captures the user-movie interactions more accurately than the matrix factorization techniques on which BMF is based.
Finally, we attempt to provide some interpretation regarding the most common (estimated) ranks. The rank of the PMF tensor is the number of states that the latent variable in the naïve Bayes model can take (cf. Section II). Since the true rank is unknown, it is difficult to determine how good the estimates are. However, the performance metrics provide at least an indication of the quality of the rank estimates. In particular, VB-PMF infers the largest rank and achieves the best performance. Similarly, the performance of SVB-PMF and SQUAREM-PMF (with model selection criteria ValErr and AIC) is comparable, and so are the inferred/selected ranks. The BIC and DNML selection criteria result in much lower ranks with an accompanying performance degradation. These observations could be explained by remarking that the rank is connected to the complexity of the learned model, i.e., how well the model can capture the underlying patterns in the data. Note that VB-PMF and SVB-PMF were initialized with relatively complex models having rank $R=50$. Both algorithms then inferred simpler models with sufficient complexity. However, VB-PMF learned a more complex model compared to SVB-PMF by processing the entire dataset in each iteration.

### v2

@@ IV-A Performance with varying number of observations
From the MSRE and mean KLD metrics, VB-PMF performs comparably to or better than the other model order estimation techniques. Here, the ‘Oracle’ ML estimate, which assumes that the true hidden state associated with each observation is known, serves as an empirical lower bound for the MSRE and the KLD. This result can be attributed to the fact that VB-PMF detects the correct rank with minimal or no misfit as the size of the dataset increases. The underfitting observed in VB-PMF for smaller values of $T$ occurs due to the Bayesian algorithm preferring simpler models (i.e., with fewer parameters) in small-sample regimes and instead seeking a closer fit to the data. Such accuracy-complexity tradeoffs are a key part of the AIC, BIC, and DNML criteria; however, with the Bayesian approach, they occur implicitly during the inference procedure, making manual model order selection unnecessary. SQ-Thresh (High) has the lowest accuracy despite using the true model order (which, for illustration purposes, was assumed to be known in advance) due to a poor fit to the data. Moreover, the runtime plots show that VB-PMF provides significant gains in computational efficiency, even though the runtime increases linearly with the dataset size $T$. This is because the rank is estimated via a single training run, while the other algorithms need to be trained afresh for each candidate rank. The computational complexity of the CTF approaches is dominated by the computation of the joint CPD from the lower-order marginals. Hence, it does not grow as rapidly with $T$ as VB-PMF. However, for larger dataset sizes, VB-PMF still exhibits an advantage over CTF in terms of accuracy.
IV-B Performance with varying ranks and outage probabilities
Next, we study the performance of VB-PMF for various ranks and outage probabilities. The ranks considered are $R=\{5,10,15\}$, the outage probabilities are $p=\{0,0.1,0.3,0.5\}$, while the number of observations is set to $T=10^{5}$. Figs. 4 and 5 presents the mean estimated rank and the mean KLD as a function of the outage probability. The MSRE plots showed a similar trend to the KLD plots and were therefore not included for brevity. For $R=5$, VB-PMF estimates the correct rank for all outage values. This also holds for $R=10$ and $p=\{0,0.1\}$. In these cases, even with missing observations, the data is sufficient to enable the algorithm to correctly learn the low-rank structure. In the remaining cases, the underfitting by VB-PMF occurs due to the accuracy-complexity tradeoff described in Section IV-A. This can be understood more clearly in Fig. 5, where VB-PMF achieves the best KLD performance in all cases except $R=5$, where SQ-BIC and SQ-DNML also show similar results. Note that SQ-BIC and SQ-DNML estimate the correct rank across all outage values for $R=5$. However, for higher rank and outage values, these two techniques prefer smaller models and heavily penalize models with more parameters, which affects their accuracy. In comparison, SQ-AIC selects larger models due to the weaker AIC penalty. A large variance in the estimated rank is observed when using validation errors as in the CTF approaches, especially for higher ranks and outage probabilities. Moreover, as the outage increases, the accuracy of the CTF methods is affected due to less reliable marginal PMF estimates, particularly for CTF5D-ValErr. Overall, we conclude that VB-PMF provides the best accuracy-complexity tradeoff while also estimating the rank without cross-validation.
Figure 6: Sensitivity of VB-PMF to the hyperparameter $\alpha_{\lambda}$ for $T=10^{5}$ and $p=0$. Left: The evidence lower bound (ELBO) averaged over various values of $\alpha_{\lambda}$. The shaded region represents one standard deviation away from the mean. Right: Mean estimated rank versus different values of $\alpha_{\lambda}$.
In Fig. 6, we investigate the effect of the choice of the hyperparameter $\alpha_{\lambda}$ on the ELBO convergence and the rank estimation performance of VB-PMF. The values considered are $\alpha_{\lambda}=\{10^{-9},10^{-8},\dotsc,10^{-1},0.5,1\}$ for $T=10^{5}$ observations and outage probability $p=0$. The ELBO figure shows the mean, taken over the values of $\alpha_{\lambda}$, and the region of one standard deviation from the mean. We observe that the convergence of the ELBO does not depend significantly on the choice of $\alpha_{\lambda}$. In addition, the estimated rank is constant for $\alpha_{\lambda}\leq 10^{-2}$ and is accurately obtained for $R=\{5,10\}$. For $R=15$, some underfitting occurs due to insufficient data, as already seen in Fig. 4. However, setting $\alpha_{\lambda}=\{10^{-1},0.5,1\}$ results in less sparse loading vector estimates, leading to rank estimates that are higher than required. A closer inspection of the estimated loading vectors in these cases revealed that although the unnecessary latent components were much smaller than the desired ones, they were not sufficiently suppressed, resulting in values greater than the pruning threshold $\alpha_{\lambda}/T$. Therefore, these components could not be removed without manual inspection or an ad hoc specification of the pruning threshold. In conclusion, it is recommended to set $\alpha_{\lambda}\ll 1$ to enable “automatic” pruning of unnecessary latent components.
V Real-Data Experiments
Figure 7: Classification accuracy and F1 score, averaged over 50 trials, on various datasets from the UCI Machine Learning Repository. Results from the top two algorithms are highlighted in bold. Our proposed method (VB-PMF) achieves results comparable to the random forest benchmark across all datasets.
In this section, we illustrate the effectiveness of VB-PMF for various statistical learning tasks using real data. Specifically, we evaluate performance in data classification and item recommendation (rating prediction) tasks. For comparison, we consider SQ-AIC, SQ-BIC, SQ-DNML, CTF3D-ValErr (see Section IV for a description), as well as some commonly used classical approaches which serve as benchmarks for each task.
V-A Data Classification
For this task, we select 5 datasets from the UCI Machine Learning Repository [38], namely, Credit Approval (Credit), Occupancy Detection (Occupancy), Congressional Voting Records (Votes), Iris, and Website Phishing (Website). The first three datasets contain 2 classes (binary classification) while the last two datasets have 3 classes (multiclass classification). In order to apply our approach (and other PMF estimation algorithms), any continuous features are discretized.
Our goal is to estimate the class label of test samples using the joint PMF of the features and the labels, estimated from training data. Given an estimated joint PMF model $\{\widehat{\bm{\lambda}},\{\widehat{\bm{A}}_{n}\}_{n=1}^{N}\}$ and a test sample $\bm{y}_{\rm test}=[y_{1},\dotsc,y_{N-1},y_{N}]$ where $y_{N}$ is the unknown class label, an estimate is found via MAP estimation
$\widehat{y}_{N}$ $=\underset{i_{N}\in\{1,\dotsc,I_{N}\}}{\arg\,\mathrm{max}}\,\mathsf{Pr}(i_{N}\,|\,y_{1},\dotsc,y_{N-1})$ (41)
$=\underset{i_{N}\in\{1,\dotsc,I_{N}\}}{\arg\,\mathrm{max}}\,\sum_{r=1}^{R}\widehat{\lambda}_{r}\widehat{\bm{A}}_{N}(i_{N},r)\prod_{n=1}^{N-1}\widehat{\bm{A}}_{n}(y_{n},r).$
For VB-PMF, an 80% / 20% train / test split is applied to each dataset. For the other PMF estimation algorithms, a 70% / 10% / 20% train/validate/test split is applied, where 10% of the data forms a validation set for rank estimation. The results are averaged over 50 independent trials, each with a different dataset split. As a benchmark, we consider the well-established random forest (RF) classifier [39] which is known to perform particularly well on tabular data. The predictive performance of the algorithms is evaluated in terms of the classification accuracy and the F1 score [40], which measures how a classification model balances the correct identification of positive instances (precision) with the minimization of false positives and false negatives (recall) [^6] For the multiclass classification experiments, we compute the F1 score based on the per-class average precision and recall. This is one of the variants of the so-called macro-F1 score [40]..
Fig. 7 shows the classification accuracy and the F1 score achieved by all approaches on the UCI datasets. Firstly, VB-PMF consistently outperforms the other PMF-based approaches across all datasets. The Bayesian model order selection mechanism integrates over parameter uncertainty rather than relying on point estimates of the parameters, leading to more robust and stable model selection. Moreover, the Bayesian formulation automatically balances model fit and complexity through the marginal likelihood, thus avoiding the need for ad hoc model selection procedures such as cross-validation or information criteria like AIC, BIC, and DNML.
Beyond comparisons within the PMF family, it is also useful to evaluate VB-PMF in relation to a widely used classifier such as the random forest (RF). VB-PMF achieves classification accuracies and F1 scores within approximately 1% of the RF benchmark across all datasets, even though it is a general-purpose probabilistic model not specifically designed for classification. These results indicate that the joint PMF model learned by VB-PMF is effective for classification tasks. In addition to its competitive predictive performance, VB-PMF offers advantages that purely discriminative models like RF do not provide. By estimating the full joint PMF of the features and labels, VB-PMF yields a richer representation of the data than discriminative models that approximate only conditional class probabilities.
Fig. 8 exemplifies some insights from the Congressional Voting Records dataset [38] using a low-rank model estimated by VB-PMF. The probability of belonging to a certain party (class) given the latent components reveals various voting blocs such as “strong Republican” (components 1, 3, and 6), “strong Democrat” (components 2 and 4), “bipartisan” (component 5), and “slightly Republican” (component 7). Furthermore, considering the probability of voting “Yes” on a certain issue (feature) given the latent components, some meaningful ideological differences within each party emerge. For instance, although blocs 1 and 3 are Republican, they differ on issues such as arms control (anti-satellite-test-ban) and foreign policy (aid-to-nicaraguan-contras). Similarly, there are clear differences between Democrat blocs 2 and 4 on crime and foreign policy. Finally, Republican bloc 6 consists of moderates who mostly take centrist positions. Such insights are not available to purely discriminative models such as RF, highlighting the added benefit of low-rank joint PMF models.
V-B Item Recommendation
Figure 8: Insights from the Congressional Voting Records dataset [38] using a low-rank model estimated by VB-PMF. The selected features are (a) physician-fee-freeze, (b) el-salvador-aid, (c) aid-to-nicaraguan-contras, (d) anti-satellite-test-ban, (e) adoption-of-the-budget-resolution, and (f) crime. The latent components uncover various voting blocs within the Democrat and Republican classes that purely discriminative models cannot reveal.
@@ V-B Item Recommendation
We evaluate the performance of the proposed algorithms in a real-world movie recommendation application. For this purpose, we utilize the MovieLens 10M dataset [41] which contains movie ratings on a 5-star scale with half-star increments, i.e., $\{0.5,\dotsc,5.0\}$. From the entire dataset, we select the ratings of $N=100$ top-rated movies [^7] It is common in collaborative filtering applications to select popular items in a preprocessing step (see, e.g., [42, 43]). Since we estimate a multivariate PMF of all movie ratings, selecting popular items can be viewed as a dimensionality reduction step. In addition, the accuracy of the estimated PMF depends on the number of data samples (i.e., users) available. Thus, it would not be feasible to model the joint PMF of all 10,000 movies with a relatively small number of users (72,000) in the dataset. Our choice to use the 100 most rated movies results in a movie-user rating matrix with about $94\%$ of the total number of users, and with reasonable dimensions such that a reliable joint PMF estimate can be obtained.. We perform some preprocessing by omitting users who have rated only one or none of the selected movies and mapping the ratings to a discrete scale $\{1,\dotsc,10\}$. Thus, we extract a movie-user rating matrix $\bm{Y}\in\mathbb{R}^{100\times T}$, where the number of users T=67,741T=67,\!741. The percentage of missing ratings in the rating matrix is about $72\%$ (outage probability $p=0.72$). Since we are dealing with real data, the true rank (model order) of the joint PMF is unknown.
@@
We aim to estimate the joint PMF (that is, $\{\bm{A}_{n}\}_{n=1}^{N}$ and $\bm{\lambda}$) of the ratings (i.e., the random variables of all 100 movies) and use the estimated PMF to predict the missing ratings. In this scenario, the joint PMF is a 100-dimensional tensor $\bm{\mathcal{X}}\in\mathbb{R}^{I_{1}\times\cdots\times I_{N}}$ with each dimension $I_{n}=10$, $n=1,\dotsc,100$. We run 50 Monte Carlo trials in which 80% of the ratings are used to estimate the joint PMF, while 20% are used for testing. In each trial, a different training/testing split is applied to the rating matrix.
@@
During the testing phase, for each test sample, one rating is hidden and predicted using the estimated PMF. The predicted rating is found by computing its conditional expectation with respect to the estimated joint PMF. In particular, let the current test sample be $\bm{y}_{\rm test}=[y_{1},\dotsc,y_{N}]^{\mathsf{T}}$ and suppose that $y_{N}=0$, i.e., hidden. The predicted rating is then
$\widehat{y}_{N}=\sum_{i_{N}=1}^{I_{N}}i_{N}\mathsf{Pr}(i_{N}\,|\,y_{1},\dotsc,y_{N-1}),$ (42)
where the conditional PMF can be computed from the joint PMF via Bayes’ theorem.
Recall that CTF3D is based on a coupled factorization of third-order marginal PMFs [7] (cf. Section IV). For this experiment, with $N=100$, computing all $N\choose 3$ third-order marginals is prohibitive in terms of computational complexity and memory requirements. Therefore, we consider a randomly chosen subset consisting of a quarter of all possible triplet combinations. It has been shown in [11] that this strategy reduces computational complexity with a relatively minor loss in accuracy.
As a benchmark, we consider biased matrix factorization (BMF) [44], a standard and widely used method for recommender systems. For BMF, the matrix rank is chosen by observing validation errors (ValErr) using 10% of the ratings as a validation set. Finally, we also calculate the user average, the movie average, and the global average, which serve as naïve predictors of the missing ratings. The quality of the predictions is evaluated by computing the RMSE and the mean absolute error (MAE) between the predicted and true ratings.
Table II presents results from the item recommendation experiment. VB-PMF and CTF3D-ValErr achieve comparable results, with the latter having a slightly better MAE performance. However, VB-PMF has a much lower execution time due to the efficient VI-based implementation. In addition, no lower-order marginals need to be computed for VB-PMF and the SQ-based approaches, which also saves computation time. Note that for a fairer runtime comparison, the candidate ranks for CTF3D and the SQ-based methods were tested simultaneously by employing parallel computation. These results further demonstrate the advantages of the Bayesian framework underpinning our approach in terms of providing a good data fit as well as efficient rank estimation without manual tuning. Finally, VB-PMF (and the other PMF-based approaches) outperforms BMF, showing the effectiveness of our approach (and low-rank PMF models in general) in recommendation tasks.
TABLE II: Performance on the MovieLens 10M dataset considering the 100 most rated movies. Since rank selection for CTF3D and SQ was done in parallel, we present the worst-case (among the candidate ranks) average runtime. The runtimes for the benchmark approaches are negligible and are not shown.
Algorithm RMSE MAE
Runtime
VB-PMF 0.872 $\pm$ 0.008 0.665 $\pm$ 0.006 72.44
CTF3D-ValErr 0.871 $\pm$ 0.008 0.659 $\pm$ 0.006 737.58
SQ-AIC 0.899 $\pm$ 0.011 0.686 $\pm$ 0.008 344.55
SQ-BIC 0.926 $\pm$ 0.008 0.708 $\pm$ 0.005 344.55
SQ-DNML 0.926 $\pm$ 0.008 0.708 $\pm$ 0.005 346.09
BMF 0.935 $\pm$ 0.018 0.706 $\pm$ 0.010
User average 0.989 $\pm$ 0.007 0.780 $\pm$ 0.005
Movie average 1.121 $\pm$ 0.006 0.902 $\pm$ 0.004
Global average 1.015 $\pm$ 0.005 0.839 $\pm$ 0.004